Presentation on theme: "14 June 2007VELOEddy Jans 0 Recent VELO-installation++ activities June 14, 2007 Potmeter calibration Bake-out Cabling & Commissioning."— Presentation transcript:
14 June 2007VELOEddy Jans 0 Recent VELO-installation++ activities June 14, 2007 Potmeter calibration Bake-out Cabling & Commissioning
14 June 2007VELOEddy Jans 1 Potmeter calibration Fred Schimmel has written PVSS-code to do the job: buitengebied: simultaneously XA: 28 2 mm XC: -28 -2 mm step , default =2 mm binnengebied: XC=-10; XA= 10 -4.5 mm, step , default =1 mm XA=10; XC= -10 4.5 mm, step , default =1 mm y-calibrationXA=2, XC=-2, Y= -4.5 4.5 -4.5 mm step , default =0.5 mm (See next talk by HdV) Limits, steps, etc. OK ??
14 June 2007VELOEddy Jans 2 Bakeout of VELO-tank & LHCb-beampipe Everything (=beam pipe, neon injection line, flanges of VELO-tank, 2 ion getter pumps) equipped with heaters, isolation and thermocouples by AT-VAC.
14 June 2007VELOEddy Jans 5 Neon injection system on balcony Bake out campaign: from 14h00 on Wednesday June 6 till 17h00 on Tuesday June 12
14 June 2007VELOEddy Jans 6 Temperature profile VELO @ 120 o C VELO @ 150 o C Test bakeout RF-boxes Bakeout RF-boxes VELO @ 100 o C VELO cooldown
14 June 2007VELOEddy Jans 7 Pressure profile VELO to 120 o C VELO to 150 o C Bakeout RF-boxes unvoluntary air injection (see plots HdV) startup of ion getter pumps VELO to 100 o C beambeam det.det. VELO cooldown 10 -10 10 -8 10 -6 10 -4
14 June 2007VELOEddy Jans 8 System is presently being vented with ultra-pure Neon. Bypass valve may NOT be opened, i.e. NO movements allowed. System will be kept “locked”, i.e. no neon supply and no balancing of pressures in beam and detector volumes. End pressures: beam@VELO: 5x10 -10, beam@MUON: 2x10 -11 VELO-detector: 3x10 -8 mbar
14 June 2007VELOEddy Jans 9 Cabling & Commissioning All HAPTAS temperature cables are connectorized by PdG + JS. Connectorizing of analog cables by INEO in D3 is ongoing, till end July. HV is almost finished LV has not started yet First crate with TELL1’s installed in D3 and configured. First crate with control boards installed on P1. Control + readout of one VELO-slice this week. Then readout of a dummy module with full DAQ-chain.
14 June 2007VELOEddy Jans 10 What’s next, apart from everything else ?? finish one detector half completely perform turn-over test with 2 cranes organize transport from Meyrin IP8 + through shaft prepare area (extra platform, store cables, clean, etc.) open VELO-tank install fitting frame measure remove fitting frame install detector half flush with dry air (how ?) and measure IV-curves or evacuate, measure IV-curves, vent with Neon finish PLC-coding motion system ELOG has moved to the following CERN Web-site http://lbweb01.cern.ch:8080
Your consent to our cookies if you continue to use this website.